# Dynamic Prompt Generation for Interactive 3D Medical Image Segmentation Training

### v2

@@ 4.3 Results on final testing set
Rank Team DSC_AUC NSD_AUC DSC_Final NSD_Final
1 aim_coreset 3.104 3.232 0.801 0.838
2 norateam_coreset 2.911 2.970 0.754 0.775
3 yiooo_coreset 2.896 2.898 0.745 0.749
4 lexor_coreset 2.501 2.572 0.625 0.643
5 ahus_coreset 2.404 2.266 0.627 0.597
6 hanglok_coreset 2.120 2.012 0.553 0.533
7 cemrg_coreset 1.885 1.624 0.476 0.410
8 sail_coreset 1.654 1.593 0.413 0.398
9 dtftech_coreset 1.591 1.079 0.417 0.284
10 owwwen_coreset 0.956 0.496 0.239 0.124
@@ 4.4 Limitation and Future Work
Our primary limitation is the dependency on bounding box initialization. While our method achieves strong performance when properly initialized, it fails for anatomical structures that lack natural bounding regionsparticularly distributed structures like vessels and airways. The central-third heuristic we employ as a fallback proves inadequate, preventing our method from handling approximately half of MRI cases where bounding boxes are unavailable.

### v3

@@ 4.3 Results on final testing set
Modality Methods DSC AUC NSD AUC DSC Final NSD Final
CT SAM-Med3D 2.0801 1.7132 0.5311 0.4385
VISTA3D 2.1644 1.9529 0.5682 0.5181
SegVol 2.3843 2.3523 0.5961 0.5881
norateam 2.8719 2.7858 0.7462 0.7290
lexor 2.3141 2.1750 0.5785 0.5437
ahus 2.3488 2.0517 0.6215 0.5518
aim 3.0628 3.0652 0.7907 0.7966
sail 1.3339 1.0638 0.3337 0.2665
yiooo 2.8254 2.6523 0.7267 0.6851
MR SAM-Med3D 2.1169 2.1185 0.5355 0.5371
VISTA3D 2.4102 2.5065 0.6327 0.6690
SegVol 2.7383 3.0322 0.6846 0.7581
norateam 3.0921 3.3433 0.7938 0.8608
lexor 2.7082 2.9463 0.6771 0.7366
ahus 2.4503 2.5190 0.6257 0.6462
aim 3.2233 3.5378 0.8292 0.9101
sail 1.9798 2.1028 0.4948 0.5256
yiooo 3.0569 3.3125 0.7819 0.8487
PET SAM-Med3D 2.3133 1.7599 0.5826 0.4465
VISTA3D 1.7576 1.2811 0.4539 0.3332
SegVol 2.7847 2.2974 0.6962 0.5744
norateam 3.2011 2.8788 0.8051 0.7265
lexor 2.5867 2.1837 0.6467 0.5459
ahus 2.2783 1.8356 0.5614 0.4565
aim 2.8622 2.4589 0.7665 0.6758
sail 2.0661 1.6516 0.5165 0.4129
yiooo 3.0282 2.6575 0.7730 0.6907
US SAM-Med3D 0.6009 0.4005 0.1502 0.1001
VISTA3D 0.7571 0.8212 0.2378 0.3116
SegVol 0.7931 1.2851 0.1983 0.3213
norateam 1.8486 2.4444 0.5294 0.7211
lexor 0.9981 1.4705 0.2495 0.3676
ahus 1.7853 2.0378 0.4744 0.5462
aim 1.8219 2.5373 0.5100 0.7335
sail 0.7624 1.1207 0.1879 0.2717
yiooo 1.5276 1.9049 0.4500 0.5894
Microscopy SAM-Med3D 0.3123 0.0311 0.0781 0.0078
VISTA3D 1.8416 2.5043 0.4701 0.6285
SegVol 2.7359 3.8765 0.6840 0.9691
norateam 3.3361 3.9719 0.8477 0.9952
lexor 2.7166 3.9274 0.6791 0.9818
ahus 2.5669 3.7553 0.6716 0.9410
aim 2.9608 3.9324 0.7884 0.9885
sail 2.3888 3.6765 0.5972 0.9191
yiooo 3.1926 3.9722 0.8092 0.9943
@@ 4.4 Limitation and Future Work
Our primary limitation is the dependency on bounding box initialization. While our method achieves strong performance when properly initialized, it fails for anatomical structures that lack natural bounding regions, particularly distributed structures like vessels and airways. The central-third heuristic we employ as a fallback proves inadequate, preventing our method from handling approximately half of MRI cases where bounding boxes are unavailable.